OpenStudy (anonymous):

solve: 3x/(2) + 1/x = (3x)/4

OpenStudy (anonymous):

Taking the lcm [3x^2 + 2]/2x = 3x /4 3x^2 + 2 = (2x)(3x)/4 3x^2 - 3x^2/2 = -2 3x^2 = -4 x^2 = -4/3 x = i sqrt (4/3), -i sqrt(4/3)